suppose you are sewing a braid border on the edges of a quilt. the quilt is 20in. longer than it is wide. you need 292in. of braid to cover the edges of the quilt. what are the dimensions of the quilt.

Answer by checkley71(8403)

L=W+20
2(W+20)+2W=292
2W+40+2W=292
4W=292-40
4W=252
W=252/4
W=63 INCHES FOR THE WIDTH.
L=63+20=83 INCHES FOR THE LENGTH.
PROOF
2*63+2*83=292
126+165=292=292